OpenStudy (anonymous):

What is the solution set for |3x + 12| + 5 = 23?

OpenStudy (anonymous):

what do you think you do first?

OpenStudy (anonymous):

subtract the 5 over

OpenStudy (anonymous):

abs(3x+12) = 18

OpenStudy (cwrw238):

now you have 2 equations to solve 3x + 12 = 18 and 3x + 12 = -18 this is because abs values is always positive) can u solve these?